Circular No. 5909

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

OJ 287 M. Kidger, Instituto de Astrofísica de Canarias; L. Takalo, Tuorla

Observatory; and the ‘OJ94 team’ report: “Recent monitoring of OJ 287 shows it to have reached its brightest level since 1984. CCD observations by Kidger with the LAC-80 telescope (Teide Observatory, Tenerife) give the following V magnitudes: Nov. 27.131 UT, 15.11; Dec. 6.090, 14.68; 6.139, 14.37; 6.176, 14.50; 7.279, 15.13. J. Mattei (AAVSO) and D. Wills (University of Texas) report the following visual magnitude estimates: Dec. 7.284, [15.3 (D. York, Abiquiu, NM); 8.225, [14.9 (York); 9.241, 14.6 (York); 9.25, 14.80 (P. Boltwood, Ontario, CCD + V filter); 10.246, 15.0 (York); 14.031, 15.1 (G. Poyner, Birmingham, England); 14.263, 15.4 (York); 15.254, 15.3 (York); 14.660, [15.4 (S. Sakuma, Kanagawa-ken, Japan). A major outburst of this object was predicted for 1994 July/August by Sillanpää et al. (1988, Ap.J. 325, 628) on the basis that the outbursts are caused by the 11.6-yr orbital period of a binary black hole. A more recent examination of the lightcurve (Kidger et al. 1992, A.Ap. 264, 32) suggests that the true period is rather shorter and gave a revised prediction for the maximum of the outburst in early 1994 March. The recent data suggest that the rise to maximum may now have initiated, or that this may be a small precursor outburst similiar to the one observed before the main outburst in 1983. Lightcurve monitoring is urgently requested at all wavelengths to establish whether or not this is the predicted outburst and to permit the observed lightcurve behavior to be contrasted with models of the predicted propagation at different frequency ranges.”

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) D. K. Yeomans and P. W. Chodas, Jet Propulsion Laboratory, report

that they closely confirm the computations by B. G. Marsden on IAUC 5906. They note that the Jupiter impact times in July 1994 are still uncertain by a few hours, but that the new times show that the impact points (defined to be the points where the cometary fragments reach the 1-bar level in Jupiter’s atmosphere), though still on Jupiter’s dark side, are significantly closer to the morning terminator than has previously been indicated. The jovigraphic latitudes at impact range from −47^{o} to −49^{o}, and the earth-Jupiter-comet angles at impact range from 95^{o} to 100^{o}.

1993 December 17 Daniel W. E. Green